OpenStudy (babynini):

Elementary question. |u+v| = ? u=<1,3> v=<-6,2>

OpenStudy (babynini):

I got <-5,5> but apparently this is not correct..I think it has to do with the | |

jimthompson5910 (jim_thompson5910):

you got u+v = <-5,5> correct but they want the length of u+v

OpenStudy (babynini):

oo how do I do that?

OpenStudy (babynini):

|v| = sqroot (a^2+b^2)

OpenStudy (babynini):

but how do I do |u+v| ?

jimthompson5910 (jim_thompson5910):

think of u+v as one vector

jimthompson5910 (jim_thompson5910):

<a,b> = <-5,5> a = -5 b = 5

jimthompson5910 (jim_thompson5910):

then you use that formula you had posted |z| = sqrt(a^2+b^2) z = u+v

OpenStudy (babynini):

would it come out to sqroot 50 ?

jimthompson5910 (jim_thompson5910):

you can simplify that

OpenStudy (babynini):

\[5\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (babynini):

Fantastic. thanks :)

jimthompson5910 (jim_thompson5910):

no problem